OpenStudy (anonymous):

Will give medal and fan... What are the explicit equation and domain for an arithmetic sequence with a first term of 6 and a second term of 2? an = 6 − 2(n − 1); all integers where n ≥ 1 an = 6 − 2(n − 1); all integers where n ≥ 0 an = 6 − 4(n − 1); all integers where n ≥ 0 an = 6 − 4(n − 1); all integers where n ≥ 1

OpenStudy (anonymous):

@hartnn @studygeek15

hartnn (hartnn):

ok, so \(a_1 = 6, a_2 =2 \) can you find the common difference 'd' from here ?

hartnn (hartnn):

then you would just use the formula \(\Large a_n = a_1+(n-1)d\) where a1 = 1st term = 6 and d =...

OpenStudy (anonymous):

I will try it (:

hartnn (hartnn):

good, common difference is difference between consecutive terms, like 2nd term - 1st term = 3rd term -2nd term =...and so on

OpenStudy (anonymous):

Is it a?

hartnn (hartnn):

d = 2nd term - 1st term = 2-6 = ... ?

OpenStudy (anonymous):

-4... so it is D then?

hartnn (hartnn):

yes! it is, because n always starts from 1 :)

OpenStudy (anonymous):

Alright! That makes sense. Thank you so much!! :D

hartnn (hartnn):

welcome ^_^